# Unifying Summary Statistic Selection for Approximate Bayesian Computation

### v3

@@ 5.3 Population genetics model
We inferred the mutation and recombination rates of a population genetics model, a problem that has been extensively studied using ABC in general and in the context of identifying summaries in particular (Joyce & Marjoram, 2008; Nunes & Balding, 2010; Blum et al., 2013). Data were generated using the coalescent approach which considers the history of a sample of haplotypes, a set of DNA variations that tend to be inherited together because they are close together on the DNA strand (Nordborg, 2019). We present the process in terms of the equivalent forward model because it is more accessible. Under the neutral Fisher-Wright model, diploid organisms (each having paired chromosomes) reproduce sexually in discrete generations without selection pressure. Haplotypes are subject to random mutations under an infinite-sites assumption, i.e. the DNA sequence is sufficiently long that the probability of multiple mutations occurring at the same site is negligible. The model also allows for recombination, i.e. the haplotype of a gamete can be a combination of parental haplotypes. We consider a finite-sites recombination model (Hudson, 1983), i.e. the strands may only cross over at specific locations during meiosis. This may seem at odds with the infinite-sites mutation model, but we can think of the haplotype as a sequence of atomic segments, each comprising many bases.
We used a dataset of $10^{6}$ simulations from the above model generously provided by Nunes & Balding (2010). The two rate parameters were drawn from a $\mathsf{Uniform}\left(0,10\right)$ prior. For each simulation, $50$ haplotypes with $5{,}001$ base pairs were generated. Seven candidate summaries comprised a uniform distractor and six expert summaries (such as the number of unique haplotypes or “the frequency of the most common haplotype” (p. 8); see Nunes & Balding (2010) for details). We split the dataset into training, validation, and test sets comprising $9.89\times 10^{5}$, $10^{4}$, and $10^{3}$ samples, respectively.
@@
We drew $1{,}000$ posterior samples for each example in the test set (comprising $\approx 0.1\%$ of the reference table as for the benchmark dataset). Kernel density estimates of the NLP are biased for bounded parameters because probability mass can “leak” out of the support (Scott, 2015). We used a reflection-based bias correction technique to estimate the NLP (Boneva et al., 1971), i.e. each approximate posterior sample $\tilde{\theta}$ is reflected at the boundaries such that both tails of the kernel contribute to the density estimate. The results are shown in panel (b) of Fig. 6 and Appendix C in the appendix. Nonlinear regression and MDN-compressed ABC not only performed comparably, but the two methods also learned very similar summaries: After standardizing and aligning the summaries using a Procrustes transformation (Schönemann, 1966), they had a pointwise MSE of $M^{2}=0.20$ ($p$-value $<10^{-3}$ under a permutation test). MDN samples performed slightly, but not statistically significantly, better in terms of EPE and RMISE. The minimum CPE method performed worst in terms of NLP because it targets highly concentrated posteriors, although not necessarily near the true value. We did not implement the two-stage method of (Nunes & Balding, 2010) due to its computational cost and similarity with posterior mean estimation approaches (Fearnhead & Prangle, 2012; Jiang et al., 2017). Similarly, we omitted approximate sufficiency (Joyce & Marjoram, 2008) because it is sensitive to the number of histogram bins chosen for density estimation and is not suitable for multidimensional parameter spaces due to the curse of dimensionality.
@@ 5.4 Growing tree model
Inferring the parameters of dynamic network models is challenging, especially when only cross-sectional data are available. Cantwell et al. (2021) developed a pseudo-marginal Markov chain Monte Carlo (PM-MCMC) algorithm to infer the parameters of a growing tree model. Under the model (Krapivsky & Redner, 2001), a tree is grown from a single isolated node. At each step, a new node $j$ is added and connected to an existing node $i$ with probability $\propto k_{i}^{\theta}$, where $k_{i}$ is the degree of node $i$, i.e. the number of connections it has. The parameter $\theta$ controls the strength of preferential attachment: The larger $\theta$ the more likely nodes are to connect to nodes that already have many connections. The likelihood is intractable because the history, i.e. the order of addition of nodes, is unknown. There are combinatorially many possible histories, and evaluating the likelihood exactly is infeasible save for very small trees. The PM-MCMC algorithm estimates the marginal likelihood by sampling a set of histories consistent with the observed graph and averaging the conditional likelihood for each history (Cantwell et al., 2021).
@@
For subset selection methods, we used one uniform distractor and four candidate summaries: The standard deviation and Gini coefficient of the degree distribution because heavy-tailed degree distributions are indicative of large $\theta$; the diameter of the network and the maximum betweenness centrality because the existence of a central hub connecting disparate parts of the tree is indicative of large $\theta$ (Newman, 2018). For nonlinear posterior mean estimation and MDN compression, we used a two-layer graph isomorphism network (GIN) where each layer comprised a two-layer MLP with eight hidden units per layer (Xu et al., 2019). We used the constant vector of ones as input features for the GIN because nodes are indistinguishable. Features were mean-pooled across the graph after applying the GIN, and training used 32 trees per mini-batch.
@@
All methods, including using candidate summaries directly, significantly reduced uncertainty about the attachment parameter. As shown in panel (c) of Fig. 6 and Appendix C in the appendix, the PM-MCMC algorithm had the lowest RMISE, but MDN-compressed ABC and direct sampling from the MDN performed best in terms of NLP. Similar to the benchmark experiment in Section 5.2, we repeated the experiment for larger trees with $n=748$ nodes. MDN compression trained on large trees performed best ($\text{NLP}=-1.67\pm 0.02$), but summaries learned on trees with $n=100$ nodes generalized to larger trees with $\text{NLP}=-1.65\pm 0.02$. The relative NLP performance of other methods remained unchanged, but PM-MCMC performance was sensitive to the size of the grown tree and degraded severely with $\text{NLP}=1.7\pm 0.2$, much larger than the prior entropy $H\left\{\pi\left(\theta\right)\right\}=0.69$. For superlinear preferential attachment, i.e. $\theta>1$, almost every new node connects to a central hub (Krapivsky & Redner, 2001). This phenomenon is particularly pronounced for larger graphs, and it is challenging to infer histories accurately which leads to poor posterior inference. In this experiment, ABC with MDN-compressed summaries outperformed the dedicated (pseudo-marginal) likelihood-based approach.
@@ 6 Discussion
We compared various methods on a benchmark problem (Section 5.2), a population genetics model (Section 5.3), and a model for growing trees (Section 5.4). Minimizing the EPE yields highly informative summaries while achieving the long-standing goal of “find[ing] methods which do not require a preliminary subjective feature selection stage” (Prangle, 2018, p. 147). But there is no free lunch: We instead have to choose a compression and density estimation architecture. Choosing appropriate architectures can improve performance, reduce the number of simulations required (Chan et al., 2018), and even allow summaries to generalize across datasets of different sizes as demonstrated in the benchmark and growing trees experiments.
Sequential methods can reduce the computational burden of likelihood-free inference (Lueckmann et al., 2017; Papamakarios & Murray, 2016; Chen et al., 2021), but we focused on learning summaries for rejection ABC for two reasons: First, we wanted to isolate the effect of summary selection without introducing confounders. We omitted regression adjustment for ABC samples (Beaumont et al., 2002) for the same reason. Second, learning global summaries allows for amortized inference because we do not need to retrain models for each example. Investigating the interaction between sequential methods and learning summaries could shed light on how different aspects of the data inform parameters in different regions of the parameter space, as illustrated in Fig. 3.
The summaries of the mixture density networks in Section 5 can have arbitrary scales which can be problematic for ABC. We standardized summaries after extraction to mitigate this problem, but metric learning approaches could further improve ABC with MDN-compressed summaries (González-Vanegas et al., 2019). Investigating the impact of model misspecification on ABC is an active area of research (Frazier et al., 2020), and comparing the robustness of different methods should be considered in future work.
Neural density estimation is a powerful tool for likelihood-free inference, “but there is no uniformly best algorithm” (Lueckmann et al., 2021, p. 1). ABC remains a compelling approach because of its theoretical properties, and it can produce high-fidelity posteriors, especially when low-dimensional but rich summaries can be extracted from complex data.
@@ Acknowledgement
We thank Matthew Nunes for sharing simulations and John Platig for feedback on the population genetics problem. This work was supported by the U.S. National Institutes of Health under Grant R01AI138901 and R35CA220523.
@@ Appendix A EPE and expected Kullback-Leibler divergence
Chen et al. (2021) made the following proposition (see p. 2); we have adapted the notation for consistency with the main text.
@@
They subsequently argued that “the sufficiency of the learned statistics is insensitive to the choice of $\pi\left(\theta\right)$” (p. 4) and that the “approach differs from [other] methods as it is globally sufficient for all $\theta$” (p. 6). As demonstrated in Section 4.5, this assertion does not hold in general. But it is true if the likelihood belongs to the exponential family and $\mathcal{T}$ includes the sufficient statistics. The proof proposed by Chen et al. (2021) does not hold because it assumes the conclusion (p. 13).
@@ Appendix C Likelihood-based inference for the benchmark problem
We used the likelihood-based inference framework Stan (Carpenter et al., 2017) to draw samples from the posterior for the benchmark problem in Section 5.2. The algorithm uses the geometry of the posterior to efficiently draw samples from it. However, complex geometries make exploring the posterior difficult, and a change of variables can be an effective means to improve sampling. For the problem at hand, we made the change of variables $\phi=\tanh\theta$ such that the likelihood becomes
@@
and we restricted $10\leq\phi<$ 1 because the posterior is symmetric and exploring one mode is sufficient. The posterior is thus
@@
where the first term is the Jacobian accounting for the change of variables. We obtained samples of $\theta$ by applying the inverse $\tanh$ transform and randomly reversing the sign with probability $0.5$. We set the target acceptance probability to $0.99$ to ensure that there are no divergent transitions which indicate numerical instabilities (Carpenter et al., 2017).
[…]
The quality of summaries has a significant impact on the fidelity of posteriors. The table reports the negative log probability loss (NLP) and root mean integrated squared error (RMISE) for combinations of methods and experiments; reported errors are standard errors. Methods that are within one standard error of the best method are highlighted in bold. See Sections 5.2, 5.3 and 5.4 for details on the benchmark, coalescent, and growing tree experiments, respectively. Benchmark Coalescent Growing tree Method NLP RMISE NLP RMISE NLP RMISE Likelihood-based $\mathbf{1.05\pm 0.01}$ $1.24\pm 0.02$ not applicable $-0.666\pm 0.028$ $\mathbf{0.165\pm 0.002}$ MDN compression $\mathbf{1.05\pm 0.01}$ $1.24\pm 0.02$ $\mathbf{3.601\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.730\pm 0.021}$ $\mathbf{0.167\pm 0.002}$ MDN $1.08\pm 0.02$ $1.24\pm 0.02$ $\mathbf{3.595\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.727\pm 0.021}$ $\mathbf{0.167\pm 0.002}$ Nonlinear regression $1.43\pm 0.02$ $1.35\pm 0.01$ $\mathbf{3.604\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.702\pm 0.021}$ $0.172\pm 0.002$ Linear regression $1.32\pm 0.02$ $1.30\pm 0.02$ $3.677\pm 0.029$ $3.98\pm 0.03$ $-0.685\pm 0.021$ $0.174\pm 0.002$ Minimum CPE $1.13\pm 0.02$ $\mathbf{1.20\pm 0.02}$ $3.734\pm 0.033$ $3.99\pm 0.03$ $\mathbf{-0.700\pm 0.022}$ $0.169\pm 0.002$ Expert summaries $1.12\pm 0.01$ $1.25\pm 0.02$ $3.685\pm 0.025$ $4.05\pm 0.03$ $-0.579\pm 0.014$ $0.208\pm 0.002$ PLS $1.15\pm 0.02$ $1.26\pm 0.02$ $3.712\pm 0.024$ $4.05\pm 0.03$ $-0.525\pm 0.013$ $0.225\pm 0.002$ Prior $1.44\pm 0.03$ $1.36\pm 0.01$ $4.621\pm 0.003$ $5.69\pm 0.03$ $0.696\pm 0.002$ $0.801\pm 0.006$

### v4

@@ 5.3 Population genetics model
We inferred the mutation and recombination rates of a population genetics model, a problem that has been extensively studied using ABC in general and in the context of identifying summaries in particular [Joyce2008, Nunes2010, Blum2013]. Data were generated using the coalescent approach which considers the history of a sample of haplotypes, a set of DNA variations that tend to be inherited together because they are close together on the DNA strand [Nordborg2019Coalescent]. We present the process in terms of the equivalent forward model because it is more accessible. Under the neutral Fisher-Wright model, diploid organisms (each having paired chromosomes) reproduce sexually in discrete generations without selection pressure. Haplotypes are subject to random mutations under an infinite-sites assumption, i.e. the DNA sequence is sufficiently long that the probability of multiple mutations occurring at the same site is negligible. The model also allows for recombination, i.e. the haplotype of a gamete can be a combination of parental haplotypes. We consider a finite-sites recombination model [Hudson1983], i.e. the strands may only cross over at specific locations during meiosis. This may seem at odds with the infinite-sites mutation model, but we can think of the haplotype as a sequence of atomic segments, each comprising many bases.
We used a dataset of $10^{6}$ simulations from the above model generously provided by Nunes2010. The two rate parameters were drawn from a $\mathsf{Uniform}\left(0,10\right)$ prior. For each simulation, $50$ haplotypes with 5,0015{,}001 base pairs were generated. Seven candidate summaries comprised a uniform distractor and six expert summaries (such as the number of unique haplotypes or “the frequency of the most common haplotype” (p. 8); see Nunes2010 for details). We split the dataset into training, validation, and test sets comprising $9.89\times 10^{5}$, $10^{4}$, and $10^{3}$ samples, respectively.
@@
We drew 1,0001{,}000 posterior samples for each example in the test set (comprising $\approx 0.1\%$ of the reference table as for the benchmark dataset). Kernel density estimates of the NLP are biased for bounded parameters because probability mass can “leak” out of the support [Scott2015MultivariateDensityEstimation]. We used a reflection-based bias correction technique to estimate the NLP [Boneva1971DensityEstimationReflection], i.e. each approximate posterior sample $\tilde{\theta}$ is reflected at the boundaries such that both tails of the kernel contribute to the density estimate. The results are shown in panel (b) of Fig. 6 and Table 2 in the appendix. Nonlinear regression and MDN-compressed ABC not only performed comparably, but the two methods also learned very similar summaries: After standardizing and aligning the summaries using a Procrustes transformation [Schoenemann1966Procrustes], they had a pointwise MSE of $M^{2}=0.20$ ($p$-value $10^{-3}$ under a permutation test). MDN samples performed slightly, but not statistically significantly, better in terms of EPE and RMISE. The minimum CPE method performed worst in terms of NLP because it targets highly concentrated posteriors, although not necessarily near the true value. We did not implement the two-stage method of [Nunes2010] due to its computational cost and similarity with posterior mean estimation approaches [Fearnhead2012, Jiang2017]. Similarly, we omitted approximate sufficiency [Joyce2008] because it is sensitive to the number of histogram bins chosen for density estimation and is not suitable for multidimensional parameter spaces due to the curse of dimensionality.
@@ 5.4 Growing tree model
Inferring the parameters of dynamic network models is challenging, especially when only cross-sectional data are available. Cantwell2021TreePosterior developed a pseudo-marginal Markov chain Monte Carlo (PM-MCMC) algorithm to infer the parameters of a growing tree model. Under the model [Krapivsky2001GrowingNetworks], a tree is grown from a single isolated node. At each step, a new node $j$ is added and connected to an existing node $i$ with probability $\propto k_{i}^{\theta}$, where $k_{i}$ is the degree of node $i$, i.e. the number of connections it has. The parameter $\theta$ controls the strength of preferential attachment: The larger $\theta$ the more likely nodes are to connect to nodes that already have many connections. The likelihood is intractable because the history, i.e. the order of addition of nodes, is unknown. There are combinatorially many possible histories, and evaluating the likelihood exactly is infeasible save for very small trees. The PM-MCMC algorithm estimates the marginal likelihood by sampling a set of histories consistent with the observed graph and averaging the conditional likelihood for each history [Cantwell2021TreePosterior].
@@
For subset selection methods, we used one uniform distractor and four candidate summaries: The standard deviation and Gini coefficient of the degree distribution because heavy-tailed degree distributions are indicative of large $\theta$; the diameter of the network and the maximum betweenness centrality because the existence of a central hub connecting disparate parts of the tree is indicative of large $\theta$ [Newman2018Networks]. For nonlinear posterior mean estimation and MDN compression, we used a two-layer graph isomorphism network (GIN) where each layer comprised a two-layer MLP with eight hidden units per layer [Xu2019GraphIsomorphismNetworks]. We used the constant vector of ones as input features for the GIN because nodes are indistinguishable. Features were mean-pooled across the graph after applying the GIN, and training used 32 trees per mini-batch.
@@
All methods, including using candidate summaries directly, significantly reduced uncertainty about the attachment parameter. As shown in panel (c) of Fig. 6 and creftbl:methods in the appendix, the PM-MCMC algorithm had the lowest RMISE, but MDN-compressed ABC and direct sampling from the MDN performed best in terms of NLP. Similar to the benchmark experiment in Section 5.2, we repeated the experiment for larger trees with $n=748$ nodes. MDN compression trained on large trees performed best ($\text{NLP}=-1.67\pm 0.02$), but summaries learned on trees with $n=100$ nodes generalized to larger trees with $\text{NLP}=-1.65\pm 0.02$. The relative NLP performance of other methods remained unchanged, but PM-MCMC performance was sensitive to the size of the grown tree and degraded severely with $\text{NLP}=1.7\pm 0.2$, much larger than the prior entropy $H\left\{\pi\left(\theta\right)\right\}=0.69$. For superlinear preferential attachment, i.e. $\theta>1$, almost every new node connects to a central hub [Krapivsky2001GrowingNetworks]. This phenomenon is particularly pronounced for larger graphs, and it is challenging to infer histories accurately which leads to poor inference. In this experiment, ABC with MDN-compressed summaries outperformed the dedicated (pseudo-marginal) likelihood-based approach.
5.5 Computational cost
Benchmark Coalescent Growing tree
Step or Method Small Large Small Large
Data generation
Training set 00:02 00:08 unknown 01:24 01:08:06
Validation set 00:02 00:02 unknown 00:03 00:48
Test set 00:02 00:02 unknown 00:03 00:47
Training
MDN 04:18 07:13 05:14 03:28 21:30
Nonlinear regression 00:36 02:55 00:59 03:22 19:54
Inference
Likelihood-based 02:29 07:43 not applicable 07:51 01:11:01
Expert summaries 00:05 00:13 00:04 00:52 25:19
MDN compression 00:03 00:05 00:02 00:07 00:20
MDN 00:02 00:02 00:03 00:02 00:03
Nonlinear regression 00:03 00:05 00:02 00:07 00:19
Linear regression 00:04 00:13 00:03 00:52 25:19
Minimum CPE 17:04 14:15 30:38 5:39 30:17
PLS 00:08 00:17 00:16 00:52 25:19
Prior 00:03 00:03 00:02 00:02 00:02
Table 1: Computational costs for data generation, training of neural compressors, and inference. Times are as hours:minutes:seconds. “MDN” refers to drawing samples directly from the learned mixture density network, “likelihood-based” refers to samples obtained with Stan for the benchmark problem and pseudo-marginal MCMC for the growing tree model, and all other methods use ABC after extracting summaries. “Small” and “Large” refer to different sample sizes for the benchmark ($n=10$ and $n=100$ samples) and growing tree ($n=100$ and $n=748$ nodes) experiments. Training times for MDN and nonlinear regression reflect a single training run. Inference times are for the complete test set of 1,000 examples. For methods using candidate summaries (expert summaries, linear regression, minimum CPE, and PLS), times include evaluating those summaries. Data generation times for the coalescent experiment are unknown as the dataset was provided by Nunes2010.
Training neural compressors, especially MDNs, is more computationally demanding than simpler linear regression or using expert summaries directly if they are cheap to evaluate. However, as shown in Table 1, the relative cost of optimizing an MDN compared with non-linear regression decreases with increasing problem complexity as the neural compressor is responsible for the majority of the computational cost. For the benchmark with a simple compressor architecture, MDN training is approximately seven times slower than nonlinear regression. For the growing tree experiment with a graph neural network compressor, the additional cost is only 8%. Training an MDN is comparable with (small benchmark) or more computationally efficient (all other experiments) than likelihood-based inference using Stan or pseudo-marginal MCMC for the growing tree experiment. Further, optimizing an MDN is a one-time expense and can extract summaries efficiently once trained. For example, computing network summaries can be costly and must be repeated for each element of the training and test sets before running ABC [Raynal2023]. This is much slower than using a graph neural network compressor in our experiments: more than 25 minutes compared with only 20 seconds. Training and applying neural compressors is also more efficient than greedy subset selection using CPE minimization because ABC needs to be run multiple times for each example to iteratively select promising summaries.
@@ 6 Discussion
We compared various methods on a benchmark problem (Section 5.2), a population genetics model (Section 5.3), and a model for growing trees (Section 5.4). Minimizing the EPE yields highly informative summaries while achieving the long-standing goal of “find[ing] methods which do not require a preliminary subjective feature selection stage” [Prangle2018, p. 147]. But there is no free lunch: We instead have to choose a compression and density estimation architecture. Choosing appropriate architectures can improve performance, reduce the number of simulations required [Chan2018], and even allow summaries to generalize across datasets of different sizes as demonstrated in the benchmark and growing trees experiments.
Sequential methods can reduce the computational burden of likelihood-free inference [Lueckmann2017, Papamakarios2016, Chen2021MutualInfo], but we focused on learning summaries for rejection ABC for two reasons: First, we wanted to isolate the effect of summary selection without introducing confounders. We omitted regression adjustment for ABC samples [Beaumont2002ABC] for the same reason. Second, learning global summaries allows for amortized inference because we do not need to retrain models for each example. Investigating the interaction between sequential methods and learning summaries could shed light on how different aspects of the data inform parameters in different regions of the parameter space, as illustrated in Fig. 3.
The summaries of the mixture density networks in Section 5 can have arbitrary scales which can be problematic for ABC. We standardized summaries after extraction to mitigate this problem, but metric learning approaches could further improve ABC with MDN-compressed summaries [GonzalezVanegas2019]. Investigating the impact of model misspecification on ABC is an active area of research [Frazier2020], and comparing the robustness of different methods should be considered in future work.
Neural density estimation is a powerful tool for likelihood-free inference, “but there is no uniformly best algorithm” [Lueckmann2021, p. 1]. ABC remains a compelling approach because of its theoretical properties, and it can produce high-fidelity posteriors, especially when low-dimensional but rich summaries can be extracted from complex data.
@@ Acknowledgement
We thank Matthew Nunes for sharing simulations and John Platig for feedback on the population genetics problem.
Statements and Declarations
Funding
This work was supported by the U.S. National Institutes of Health under Grant R01AI138901 and R35CA220523.
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.
Code and Data Availability
The population genetics simulations were generously provided by Matthew Nunes and are available at https://github.com/onnela-lab/coaloracle. All code and other simulations are available at https://github.com/onnela-lab/summaries.
@@ Appendix A EPE and expected Kullback-Leibler divergence
Chen2021MutualInfo made the following proposition (see p. 2); we have adapted the notation for consistency with the main text.
@@
They subsequently argued that “the sufficiency of the learned statistics is insensitive to the choice of $\pi\left(\theta\right)$” (p. 4) and that the “approach differs from [other] methods as it is globally sufficient for all $\theta$” (p. 6). As demonstrated in Section 4.5, this assertion does not hold in general. But it is true if the likelihood belongs to the exponential family and $\mathcal{T}$ includes the sufficient statistics. The proof proposed by Chen2021MutualInfo does not hold because it assumes the conclusion (p. 13).
@@ Appendix C Likelihood-based inference for the benchmark problem
We used the likelihood-based inference framework Stan [Carpenter2017] to draw samples from the true posterior for the benchmark problem in Section 5.2. The algorithm uses the geometry of the posterior to efficiently draw samples from it. However, complex geometries make exploring the posterior difficult, and a change of variables can be an effective means to improve sampling. For the problem at hand, we made the change of variables $\phi=\tanh\theta$ such that the likelihood becomes
@@
and we restricted $0\leq\phi<1$ because the posterior is symmetric and exploring one mode is sufficient. The posterior is thus
@@
where the first term is the Jacobian accounting for the change of variables. We obtained samples of $\theta$ by applying the inverse $\tanh$ transform and randomly reversing the sign with probability $0.5$. We set the target acceptance probability to $0.99$ to ensure that there are no divergent transitions which indicate numerical instabilities [Carpenter2017].
Benchmark Coalescent Growing tree
Method NLP RMISE NLP RMISE NLP RMISE
Likelihood-based $\mathbf{1.05\pm 0.01}$ $1.24\pm 0.02$ not applicable $0.666\pm 0.028$ $\mathbf{0.165\pm 0.002}$
Expert summaries $1.12\pm 0.01$ $1.25\pm 0.02$ $3.685\pm 0.025$ $4.05\pm 0.03$ $0.579\pm 0.014$ $0.208\pm 0.002$
MDN compression $\mathbf{1.05\pm 0.01}$ $1.24\pm 0.02$ $\mathbf{3.601\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.730\pm 0.021}$ $\mathbf{0.167\pm 0.002}$
MDN $1.08\pm 0.02$ $1.24\pm 0.02$ $\mathbf{3.595\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.727\pm 0.021}$ $\mathbf{0.167\pm 0.002}$
Nonlinear regression $1.43\pm 0.02$ $1.35\pm 0.01$ $\mathbf{3.604\pm 0.028}$ $\mathbf{3.82\pm 0.03}$ $\mathbf{-0.702\pm 0.021}$ $0.172\pm 0.002$
Linear regression $1.32\pm 0.02$ $1.30\pm 0.02$ $3.677\pm 0.029$ $3.98\pm 0.03$ $0.685\pm 0.021$ $0.174\pm 0.002$
Minimum CPE $1.13\pm 0.02$ $\mathbf{1.20\pm 0.02}$ $3.734\pm 0.033$ $3.99\pm 0.03$ $\mathbf{-0.700\pm 0.022}$ $0.169\pm 0.002$
PLS $1.15\pm 0.02$ $1.26\pm 0.02$ $3.712\pm 0.024$ $4.05\pm 0.03$ $0.525\pm 0.013$ $0.225\pm 0.002$
Prior $1.44\pm 0.03$ $1.36\pm 0.01$ $4.621\pm 0.003$ $5.69\pm 0.03$ $0.696\pm 0.002$ $0.801\pm 0.006$
Table 2: The quality of summaries has a significant impact on the fidelity of posteriors. The table reports the negative log probability loss (NLP) and root mean integrated squared error (RMISE) for combinations of methods and experiments. “MDN” refers to directly sampling from the learned mixture density network, “likelihood-based” refers to samples obtained with Stan for the benchmark problem, and all other methods use ABC after extracting summaries. Reported errors are standard errors, and methods that are within one standard error of the best method are highlighted in bold. See Sections 5.2, 5.3 and 5.4 for details on the benchmark, coalescent, and growing tree experiments, respectively.